OpenStudy (anonymous):

Order the following from least to greatest... 0.2(Repeating),1/5,0.02 1 1/10, 1.1 (repeating), 1.101 6/5, 1 5/6, 1.3 (repeating)

OpenStudy (anonymous):

\[0.2 (Repeating), 1/5, 0.02\] 1 1/10, 1.1 (repeating), 1.101 6/5, 1 5/6, 1.3 (repeating)

OpenStudy (anonymous):

By "(repeating)" I mean it is a repeating decimal...

OpenStudy (anonymous):

1.) 0.02, 1/5, 0.2(repeating) 2.) 11/10, 1.101, 1.1(repeating) 3.) 6/5, 1.3(repeating), 1 5/6

OpenStudy (anonymous):

Sorry it took so long haha

OpenStudy (anonymous):

It's fine lol! Thanks for the help :)

OpenStudy (anonymous):

1st- 0.02, 1/5, 0.2222 2nd-1.101 1.1/10 1.1111 3rd-6/5 1.3 1.5/6

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Last one... LOL \[4.3(repeating), \frac{ 9 }{ 2 }, 4\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

Ryan the second one is wrong because its 11/10 = 1.100 1.101 1.111

OpenStudy (anonymous):

\[4.3(repeating), 4 \frac{ 3 }{ 7 }, \frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

I promise... LAST ONE! A group of gymnasts were asked to name their favorite piece of equipment. 0.33 of the gymnasts chose the vault, 4/9 chose the beam, and 1/7 chose the uneven parallel bars. List their preference in greatest to least..

OpenStudy (anonymous):

lolz it's fine :P

OpenStudy (anonymous):

Had a sub today and the lesson notes read only 1-21 but she didn't tell us that didn't want to take a chance so did them all..

OpenStudy (anonymous):

4/9, 0.33, 1/7

OpenStudy (anonymous):

Yeah, subs can make things pretty difficult at times